Jan 14, 2016 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Update BCAL clusterizer to improve agreement between simulated and generated MC (Mark)
  2. Modify covariance matrix for FCAL and BCAL to correspond to the current resolution of the detectors (Matt).
  3. Change F250 simulation code to match firmware and add flexibility for new algorithms (Mike)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Action Items: Mike looking at 125 stuff trying to sort them out. Missing things from per-event-record; will talk to David.
  3. Collaboration Meeting February 18-20, 2016.
  4. Run Preparations
    1. FCAL Update
    2. BCAL Update (Will's pi0s)
  5. Reconstruction and Simulation
    1. Eta Prime Update
  6. Calibration
    1. Calibration Challenge
  7. Any other business


Attendees: Matt (IU), Will, Mike (CMU), Justin, Adesh, John, Mark (JLab), Christina, George (Athens), Andrei, Tegan, Ahmed, Noemi, Zisis (Regina)"

  1. Announcements:
    1. Beam Coordination Meeting: the group discussed Alexandre's Wiki page for the spring 2016 run. Calorimetry was not a big topic of discussion; it is considered to be under control and is not a high priority commissioning item for spring 2016. Discussion centred on beam, magnets, trigger. These meetings will continue on a weekly bases. Exact time of meeting? Matt and sometimes Zisis have a conflict. Zisis will bring this up with Alexandre.
  2. Action Items
  3. Collaboration Meeting February 18-20, 2016. Matt: standard 30 min FCAL talk is ok and Adesh will give. Liping will give a 20 min talk. BCAL has two 30 min slots. So we are still with the 90 min of calorimeters, and this typically works for us. Zisis will assign some placeholder names for now, which we can fine tune later.
  4. Run Preparations
    1. FCAL Update:
      1. main priority is get the bases going. 15 bases need to be replaced; 10 fail to HV=0, 5 have 3% HV set below set point (sporadic short). Some bases have 15mA or higher current. Make improvements (get rid of bad ones and ones that generate alarms) before hall closes on Feb 4. However, replacements can also occur during accelerator beam studies on Wednesdays (controlled access). FCAL is on power now. No single channel has failed in 3 days. Adesh will post plots on logbook showing pi0.
    2. LED system for FCAL: was taken data before fall run. Took one run in fall. The LED will be useful for gain balancing (fixed gain adjustment); shows how much to increase/decrease pulse integral. PMT response vs HV is not as good as LED for this job (bases have changed, for example). LED pulses are stable.
    3. BCAL Update: ditto on LED. We plan to run and look at that data.
      1. (Will's pi0s): divided pi0 by total time of runs to get an estimate of rates. Will would like 3-5 times more stats than spring 2015 run to do positions, energies, z-ranges binning (e.g. 10 bins in z). Elton's estimate is 81 hours of beam time with the same rates as spring 2015.
      2. George: nothing to report.
      3. Noemi: t-w; now using pulse peak instead of pulse integral, has increased range of fits, tries to extend 3-param fit to include a linear term in ADC.
      4. Ahmed: efficiencies: wrapping up report with 0-26 sampling tried on mod 8 runs from spring 2015. This change (from 5-55) does not seem to affect the efficiency.
      5. Andrei: still working on extracting the calibration numbers from the Dec 2015 cosmic runs.
      6. Will/Muons: Andrei's ROOT trees look good and will process it.
      7. Mark: will take cosmics over the weekend at the new temperature. Look at trigger separation. It could be mod 8 data. Andrei's script uses mod 8. Rate is low enough to take mod 8.
  5. Reconstruction and Simulation
    1. Eta Prime Update: eta-prime->eta pi+ pi- and eta->2g. Vertex-R was not used; will be looked at. MC and all 6 runs from spring data. Missing eta-prime resolution (off proton) is better than the reconstructed eta-prime resolution (on proton). 1st plot, right panel, proton cut can be improved to include more protons (signal follows proton curve; dE/dx cut cleans up this plot). Use timing info to clean things up; trying to understand timing issues between MC and data. Distinguish bggen signal events (use flag); show plots for only signal and also for everything. Cuts were defined using signal MC.
  6. Calibration
    1. Calibration Challenge: nothing to report.
  7. Any other business